# Artificial Intelligence LLM Large Language Model

### v2

@@ 5.3.1. Theoretical Implications
Calibrated user perceptions and the role of cues
Our study adds to the mixed findings on calibrated trust (Wischnewski et al., 2023), by showing that customizations of an LLMaaS chatbot influence users’ trust significantly, even when their knowledge of actual system differences is limited. Furthermore, we extended the concept of calibration to LLM hallucinations and privacy, highlighting the significance of calibrated AI perceptions and use in other aspects. We also extended the TrAM framework (Schlicker et al., 2025) by applying it to users’ perceived privacy and self-reported experienced hallucinations. Our results indicate that customizations of LLMaaS chatbots are perceived by users in the form of cues and can have a significant impact on their perceived privacy, trust in the system and experienced hallucinations. These findings suggest that the TrAM is particularly useful for identifying points in the user experience where misalignments between users’ perceptions and the system’s actual characteristics may occur. This makes it a valuable framework for scholars and practitioners to use when identifying or refining existing cues, or when introducing new ones, in order to foster more calibrated user perceptions.
Our findings also open up important avenues for future research on the cues underlying user judgments of AI systems. While we showed that system customizations influence user perceptions of trust, privacy, and hallucinations, it remains an open question which specific cues users rely on most when forming such judgments. A key next step for theory development in the area of calibrated AI use is to investigate what kinds of cues users notice, which ones they consider most relevant, and how well these cues reflect the actual characteristics of the system. Future studies should seek to systematically compare different cues in terms of their influence on user perceptions and their alignment with actual system properties in order to better understand how (mis)calibrations emerge in practice.
@@
This work stimulates new directions for research on the users’ detection of LLM hallucinations, particularly regarding usage context and task characteristics. Prior work has mainly focused on individual-level factors (Dang and Nguyen, 2025), explanation formats (Reinhard et al., 2025), and information characteristics (Tao et al., 2024), but has paid less attention to how system use itself shapes hallucination detection. Our results point to notable differences in the broader usage context of the two systems. Although participants were instructed to answer the questionnaire items with regard to their work- or study-related use, the reported overall usage patterns diverged substantially: within the subsample of both systems users, the customized LLMaaS chatbot was primarily used for academic or professional tasks, whereas ChatGPT use was more evenly distributed between private and academic contexts (see Section 4.4.4). It remains an open question whether such contextual differences affect users’ expectations, vigilance, or standards for verifying outputs, and consequently their likelihood of noticing hallucinations. In addition to context, future research should also consider how different usage purposes influence hallucination alertness and detection. Beyond broad categories such as information search or text creation, more nuanced dimensions of purpose, such as task complexity, topical domain, or the type of user goal (e. g., idea generation vs. factual information retrieval), might influence how carefully outputs are evaluated. Addressing these aspects through controlled experiments could enrich existing research and provide a more comprehensive understanding of hallucination detection.
@@ 5.3.2. Practical Implications
Most importantly, we advise institutions deploying LLMaaS chatbots to pay attention to the potential discrepancy between the actual and perceived trustworthiness of the system, i. e., ensuring that user perceptions are calibrated correctly in relation to the characteristics of the system. In particular, developers and deployers should ensure that surface-level cues do not create a false impression of the system’s capabilities or privacy safeguards. This could involve actively communicating residual risks, providing context-specific warnings, and offering transparency tools to foster better calibration. The results of our study showed that although a system might not hallucinate less, it can be perceived to do so. Therefore, we strongly suggest implementing a hallucination warning in the user interface. Previous research has shown that warnings can increase hallucination detection without significantly affecting perceived truthfulness of genuine content (Nahar et al., 2024). Similarly, Bo et al. (2025) found that adding a persistent disclaimer below the LLM output was the most effective at improving users’ appropriate reliance in a complex problem-solving task (“Remember to verify this information and adapt it to your specific problem as I may not have complete information about your context.”). Building on this, additional warnings could be provided when the system offers links, encouraging users to verify their accuracy and relevance to their request before using them for further purposes, e. g., in academic writing. Moreover, confidence scores presented alongside generated output and links connecting system responses to the source documents used to generate them could be effective transparency tools to support calibrated user perceptions (Leiser et al., 2023; Reinhard et al., 2025).
@@
At an institutional level, measures could include a mandatory online course providing information on the risks, limitations, and particularities of the system, which users would have to complete to activate their account for using LLMaaS chatbots. One helpful customization that could support the calibrated use of a university’s LLMaaS chatbot would be to integrate a dedicated “audit-assist” mode with structured, example-driven scaffolding features that guide users through systematic audits of model behavior. Such scaffolding, as shown in a recent study (Prabhudesai et al., 2025), can improve students’ critical engagement with LLMs, including their ability to identify biases. Furthermore, alerts could be added to the interface when users ask for medical, legal or financial advice, or when they enter personal data (e. g., names, email addresses, etc.), to facilitate critical output evaluation in high-risk contexts. To encourage appropriate user behavior particularly in relation to privacy, one option for developers of LLMaaS chatbots would be to integrate user-led data minimization tools that help users detect and remove personal information from their prompts (Zhou et al., 2025).
Another promising institutional measure is the integration of usage analytics and feedback loops. By collecting anonymized data on chatbot interactions (e. g., hallucination reports, privacy concerns, trust difficulties) and combining them with periodic user feedback, universities can continuously monitor risks and adapt training or interface features accordingly. Such human-in-the-loop monitoring can help calibrate user perceptions with actual system performance, ensuring that trust remains both appropriate and evidence-based.
Supporting sustainable AI usage
Moving on, organizations could also use customizations to support sustainable AI usage. Although our within-subject analysis did not support H6 regarding sustainability-aware AI use, the significant between-subject difference suggests that customization features, such as displaying the token contingent in the interface, may encourage more sustainable usage behavior. Institutions might leverage such design choices to align AI usage with broader sustainability goals. The token contingent displayed in the chatbot’s user interface is one promising option, while further customizations could include intuitive labels indicating resource efficiency or environmental impact of different models, such as care-label-inspired badges with color-coded usage indicators for energy consumption and runtime (Fischer et al., 2024, 2025).
User acceptance as a baseline
For decision makers in organizations, our findings indicate that introducing customized LLMaaS chatbots can build on a solid baseline of user acceptance. Some customizations may even improve perceived ease of use compared to established commercial LLM chatbots. The exploratory analyses showed that, while users found both systems equally useful, they indicated higher ease of use for the customized LLMaaS chatbot compared to ChatGPT. These findings are valuable in terms of the applicability of customized LLMaaS chatbots in real-world institutional contexts, given that these two variables are key factors in predicting user acceptance of technologies (Davis, 1989), a link that has also been shown for AI systems (Choung et al., 2022; Kelly et al., 2023).
@@ 5.4. Limitations
The study has limitations that need to be taken into account in the interpretation of the results and suggest avenues for future research.
First, our main analyses were based on a within-subjects subsample of users who had experience with both systems. While this group enables controlled comparisons, it may also introduce a self-selection bias. Although key effects (e. g., privacy and hallucinations) were replicated in the between-subjects analysis, sustainability-aware AI usage was only significant in the between-subjects comparison. Future studies should therefore examine the generalizability of our findings across user groups (e. g., students vs. employees; single-system vs. multi-system users).
@@
Second, we did not control for potential order effects in the questionnaire. Due to the limited configuration options on the university’s internal survey platform, the university’s LLMaaS chatbot was always evaluated before ChatGPT. While this ordering was consistent for all participants and unlikely to explain system-specific effects alone, it remains a potential confound. Future studies should employ counterbalanced survey designs to rule out such effects.
@@
Third, our study relied exclusively on self-report measures. Future studies should incorporate behavioral data, such as prompt input, output verification actions, or system usage logs, to complement self-reports. Furthermore, the internal consistency of our newly developed scale measuring cautious behaviour towards hallucinations was low. This may be due to heterogeneity in the construct, as it includes both proactive (e. g., adjusting prompts) and reactive (e. g., checking outputs) strategies. Future research should consider splitting and extending this scale to better distinguish between these dimensions and improve psychometric quality.
@@
Fourth, although we conducted a priori power analyses, our within-subjects sample (n = 116) fell short of the planned size (n = 250). While sufficient for detecting medium effects, smaller effects, such as differences in cautious behavior or sustainability-aware AI use, may have gone undetected.
@@
Fifth, the study design does not allow for causal interpretations. While one key strength of our study is that we surveyed real users in a naturalistic setting, thereby generating findings with high ecological validity, this comes at a cost of lower internal validity. This could be addressed in future research through the use of controlled experimental designs to isolate the effects of individual features, while longitudinal studies may shed light on how user perceptions evolve over time. Such designs could also examine potential causal links between specific user perceptions, e. g., user trust and experienced hallucinations.
@@
Finally, we see potential for qualitative follow-up research to better understand how and why the patterns observed in our quantitative findings emerge. Several open-ended comments from participants hint at meaningful underlying reasoning and motivations, e. g.: “I always use [the university chatbot] when I have concerns about data privacy regarding the content submitted to ChatGPT.” Future studies could conduct in-depth interviews with users to explore the use of cues in system assessments and decision-making strategies when choosing between different LLM chatbots.
$5.$ Conclusion
We examined how user-salient customizations in LLMaaS chatbots, such as institutional branding and interface changes, shape user perceptions and behavior. Participants reported greater trust, fewer perceived hallucinations, and lower privacy concerns when using the university’s customized chatbot compared to ChatGPT – despite both systems being based on the same underlying LLM technology. These differences highlight the influence of contextual and interface-level cues on users’ system assessments. Our findings emphasize the importance of calibrated user perceptions not only with regard to trustworthiness, but also in relation to privacy and hallucination risk. In doing so, our work contributes to the TrAM and the concept of calibrated trust, expanding both to include more facets of human-AI interaction. Practically, we provide evidence-based guidance on AI implementation. The careful design of interfaces, features, and integration strategies for LLMaaS chatbots can promote their appropriate use in line with their risks and limitations.
We encourage future research to explore which specific cues users rely on when forming their perceptions and how such alignments affect downstream behavior. Our work therefore opens new research directions to investigate how customization aspects impact users.
@@ Acknowledgements.
We would like to thank Markus Pauly, Nico Föge, and Daniel Klippert for their valuable support with the statistical analyses. We are also grateful to Lin Kyi, Gabriel Lima, and Yixin Zou for their helpful feedback. Our sincere thanks go to all participants for contributing their time and insights, as well as to the involved university stakeholders for their continuous support in enabling this study. This work was supported by the Sponsor Research Center Trustworthy Data Science and Security https://rc-trust.ai (https://rc-trust.ai), one of the Research Alliance Centers within the UA Ruhr (https://uaruhr.de).
@@ Appendix A Demographics
Figure A1. Percentage of participants using different chatbot systems.:
@@ Appendix B Correlation Tables
Note. $\alpha$ = Cronbach’s Alpha. Values in brackets = 95% CI. * $p<.05$, ** $p<.01$.
@@
Note. $\alpha$ = Cronbach’s Alpha. Values in brackets = 95% CI. * $p<.05$, ** $p<.01$.
@@ Appendix C Descriptive Results
Figure C1. Usage frequency of the university’s customized LLMaaS chatbot vs. ChatGPT in the subsample of both systems users ($n=116$).:
Figure C2. Purpose of use for the university’s customized LLMaaS chatbot vs. ChatGPT in the subsample of both systems users ($n=116$).:
@@ Appendix D Controlling for Confounding Variables
[htb] Paired t-test comparisons of the customized LLMaaS chatbot (UniGPT) vs. ChatGPT – Non-switchers ($n=68$) Dependent Variable M (UniGPT) SD (UniGPT) M (ChatGPT) SD (ChatGPT) t df p Holm $\alpha$1 d2 95% CI3 Privacy concerns (H5) 2.32 0.89 3.46 0.80 -9.34 67 ¡ 0.001 0.00833 -1.13 [-1.44, -0.82] Experienced hallucinations (H4) 3.10 1.05 3.86 0.84 -6.41 67 ¡ 0.001 0.01 -0.78 [-1.05, -0.50] Trust (H1) 3.60 0.71 3.09 0.81 5.97 67 ¡ 0.001 0.0125 0.72 [0.45, 0.99] Sustainability-aware AI use (H6) 2.30 1.13 2.25 1.16 0.91 67 0.37 0.0167 0.11 [-0.13, 0.35] Cautious behavior toward hallucinations (H3) 4.20 0.67 4.19 0.66 0.10 67 0.92 n/a4 0.01 [-0.23, 0.25] Usefulness5 4.03 0.74 3.97 0.80 0.68 67 0.50 0.08 [-0.16, 0.32] Ease of use5 4.52 0.67 4.46 0.58 1.02 67 0.31 0.12 [-0.11, 0.36]
  • 1
    Holm-adjusted significance threshold per hypothesis rank (sequential correction).
  • 2
    Cohen’s d is reported as the standardized effect size for each comparison.
  • 3
    Confidence intervals (CI) for Cohen’s d are not adjusted for multiple comparisons.
  • 4
    Holm-adjusted thresholds are only assigned while p-values remain below $\alpha$; correction stops after the first non-significant test.
  • 5
    Ease of Use and Usefulness were analyzed exploratively and were not part of the Holm correction.
@@
ExperiencedHallucinations: System $0.537^{***}$ $0.154$ 3.49-3.49 <0.001<0.001
ExperiencedHallucinations: System × Knowledge -0.027 $0.046$ 0.60-0.60 $0.553$
@@
Appendix E Exploratory Analyses

### v3

@@ 5.2.1. Theoretical Implications
We draw on the TrAM model (Schlicker et al., 2025) to theorize why there might be differences in perceptions between customized LLMaaS and commercially available alternatives. Specifically, we theorize that perceived system characteristics only partially link to actual system characteristics and that there are cues experienced by users that lead to differences in perceived system characteristics that are unrelated to the actual system characteristics. We refer to the concept of calibration to describe the resulting fit between actual system characteristics and perceived system characteristics, a concept that has previously been considered primarily in relation to trust (Parasuraman and Riley, 1997; Lee and See, 2004).
In general, we see that users displayed higher trust, perceived greater privacy and perceived less hallucinations with the customized LLMaaS chatbot. At first glance, this suggests that users are able to calibrate their perceptions towards an LLM chatbot, as, e. g., the university’s customized chatbot has stronger privacy safeguards and users have less concerns. However, we question whether the extent of the perceived difference is justified and if users’ consequent behavior is “right”. First, we question how well users understand the differences between the systems, given the results of the knowledge test (see Section 4.3). Second, an appropriate calibrated perception is further questioned by the results of the hallucination benchmark evaluations, which allowed us to relate users’ subjective experiences to objective performance metrics. Participants perceived fewer hallucinations with the customized LLMaaS, although the TruthfulQA benchmark suggests that this system may actually be more prone to hallucinations. Accordingly, the anticipated link between the system’s lower temperature setting and reduced hallucination tendency (see Figure 1) did not hold. In sum, this divergence highlights a critical gap in users’ hallucination alertness calibration, that is, their ability to align vigilance against hallucinations with the system’s actual risk. Users may underestimate the risk of hallucinations due to trust-inducing cues such as branding, institutional context, and lack of persistent warnings. When this finding is considered alongside the significantly higher perceived privacy in LLMaaS, the question arises as to whether users are overtrusting the customized LLMaaS chatbot.
Taken together, the findings suggest that the TrAM is useful for identifying points in the user experience where misalignments between users’ perceptions and the system’s actual characteristics may occur. This makes it a valuable framework for scholars and practitioners to use when identifying or refining existing cues, or when introducing new ones, in order to foster more calibrated user perceptions.
Toward a Causal Understanding of User Perception Differences.
Our findings also open up avenues for future research on the cues underlying user judgments of AI systems. While we showed that user perceptions differed and we offer theoretically grounded interpretations as to how system customizations might have contributed to the results, the causality of the findings, e.g., with regards to the question of which specific cues users rely on most when forming such judgments, remains an open question. A key next step is to experimentally and systematically compare customization cues in terms of their influence on user perceptions and their alignment with actual system properties to better understand how (mis)calibrations emerge in practice. Future studies might incorporate experimental lab settings to test mediation analysis to model the causal relations between system type as experimental conditions, customization choices as mediators and user perceptions as outcomes.
@@
This work stimulates new directions for research on users’ detection of LLM hallucinations, particularly regarding usage context and task characteristics. Prior work has mainly focused on individual-level factors (Dang and Nguyen, 2025), explanation formats (Reinhard et al., 2025), and information characteristics (Tao et al., 2024), but has paid less attention to how system use itself shapes hallucination detection. Our exploratory results point to notable differences in the broader usage context of the two systems (see Section 4.5.4). It remains an open question whether such contextual differences affect users’ expectations, vigilance, or standards for verifying outputs. Future research should also examine how finer-grained usage purposes – such as task complexity, topical domain, or goal type (e. g., idea generation vs. factual retrieval) – shape users’ scrutiny of outputs, providing a more comprehensive understanding of hallucination detection.
@@ 5.2.2. Practical Implications
We advise institutions deploying LLMaaS chatbots to pay attention to the potential discrepancy between the actual and perceived trustworthiness of the system. In particular, developers and deployers should ensure that surface-level cues do not create a false impression of the system’s capabilities or privacy safeguards. This could involve actively communicating residual risks, providing context-specific warnings, and offering transparency tools. The results of our study showed that although a system might not hallucinate less, it can be perceived to do so. Therefore, we suggest implementing a hallucination warning in the user interface. Previous research has shown that warnings can increase hallucination detection without significantly affecting perceived truthfulness of genuine content (Nahar et al., 2024). Similarly, Bo et al. (2025) found that adding a persistent disclaimer was the most effective at improving users’ appropriate reliance in a complex problem-solving task. Building on this, additional warnings could be provided when the system offers links, encouraging users to verify their accuracy and relevance. Moreover, confidence scores and links connecting system responses to the source documents used to generate them could be effective transparency tools for supporting users’ trustworthiness assessment process  (Leiser et al., 2023; Reinhard et al., 2025).
@@
At an institutional level, measures could include a mandatory online course providing information on the risks, limitations, and particularities of the system, which users would have to complete to activate their account for using LLMaaS chatbots. One helpful customization to support users’ trustworthiness assessment process would be to integrate a dedicated “audit-assist” mode with structured, example-driven scaffolding features that guide users through systematic audits of model behavior. This can improve students’ critical engagement with LLMs (Prabhudesai et al., 2025). Furthermore, alerts could be added to the interface when users ask for medical, legal or financial advice, or when they enter personal data, to facilitate output evaluation in high-risk contexts. To encourage appropriate user behavior particularly in relation to privacy, one option for developers of LLMaaS chatbots would be to integrate user-led data minimization tools that help users detect and remove personal information from their prompts (Zhou et al., 2025).
@@ 5.3. Limitations
First, our main analyses were based on a within-subjects subsample of users who had experience with both systems. While this enables controlled comparisons, it may also introduce a self-selection bias. Although key effects (e. g., privacy and hallucinations) were replicated in the between-subjects analysis, sustainability-aware AI usage was only significant in the between-subjects comparison. Future studies should therefore examine the generalizability of our findings across user groups (e. g., students vs. employees; single-system vs. multi-system users). The generalizability of the results must also be discussed with regards to other (a) universities, (b) professional contexts, and (c) nationalities. In principle, case studies are transferable to the same extent as large-scale studies, when taking into account how similar the socio-organizational structures in another setting are (Tsang, 2014). As many German universities currently introduce LLMaaS (without fine-tuning models), the results of this study are likely representative, if similar customization choices are made. It is also conceivable that similar effects are observable in other knowledge-work intense professions. However, results might not transfer to other domains e.g., healthcare, where for example reliability and accuracy of systems are large concerns (Chen et al., 2025). Additionally, Germany is characterized by strict data protection rules. Universities/ organization need to act in accordance with the General Data Protection Regulation (GDPR), which students and staff are likely aware of. This limits generalization to other (especially non-European) countries, particularly with regard to the perceived privacy of systems.
@@
Second, we did not control for potential order effects in the questionnaire. Due to the limited configuration options on the university’s internal survey platform, the university’s LLMaaS chatbot was always evaluated before ChatGPT. While this ordering was consistent for all participants and unlikely to explain system-specific effects alone, it remains a potential confound.
Third, our study relied exclusively on self-report measures and exploratory benchmark evaluations. Future studies should incorporate behavioral data, such as prompt input, output verification actions, or system usage logs, to complement self-reports. Moreover, perceptual measures alone may not fully capture how customization relates to actual model behavior. Although we included two exploratory objective benchmark evaluations to assess hallucination tendencies and detection performance (see Section 4.5.1), these analyses were not part of the primary inferential design. Combining self-report measures with objective task-based performance metrics would help distinguish between effects on user perception driven by interface cues and those reflecting underlying system performance.
@@
Furthermore, the internal consistency of the scale measuring cautious behaviour towards hallucinations was low. This may be due to heterogeneity in the construct, as it includes both proactive (e. g., adjusting prompts) and reactive (e. g., checking outputs) strategies. Future research should consider splitting and extending this scale to better to distinguish dimensions and improve psychometric quality.
@@
Fourth, our within-subjects sample (n = 116) fell short of the planned size (n = 250). While sufficient for detecting medium effects, smaller effects, such as differences in cautious behavior or sustainability-aware AI use, may have gone undetected.
@@
Fifth, the study design does not allow for causal interpretations. While one key strength of our study is that we surveyed real users in a naturalistic setting, thereby generating findings with high ecological validity, this comes at a cost of lower internal validity. This could be addressed in future research through the use of controlled experimental designs isolating effects of individual features, while longitudinal studies may shed light on how user perceptions evolve over time. Such designs could also examine causal links between specific user perceptions, e. g., user trust and perceived hallucinations.
@@
5.4. Conclusion
We examined how customized LLMaaS chatbots are perceived and used differently compared to commercially available alternatives. Participants reported greater trust, fewer perceived hallucinations, and lower privacy concerns when using the university’s customized chatbot compared to ChatGPT – despite both systems being based on the same underlying LLM technology. These differences hint at the potential influence of contextual and interface-level cues on users’ system assessments. Our findings emphasize the importance of calibrated user perceptions not only with regard to trustworthiness, but also in relation to privacy and hallucination risk. In doing so, our work contributes to the TrAM and the concept of calibrated trust, expanding both to include more facets of human-AI interaction. Practically, we provide evidence-based guidance on AI implementation. The careful design of interfaces, features, and integration strategies for LLMaaS chatbots can promote their appropriate use in line with their risks and limitations.
We encourage future research to build on the results and to explore which specific cues users rely on when forming their perceptions and how such alignments affect downstream behavior in experimental settings to complement the findings of this work with causal data. Our work, therefore, opens new research directions to investigate how customization aspects impact users.
@@ Acknowledgements.
We thank Markus Pauly, Nico Föge, and Daniel Klippert for their support with the statistical analyses. We are grateful to Lin Kyi, Gabriel Lima, and Yixin Zou for their thoughtful feedback on the manuscript. We also acknowledge the support of the involved parties at Ruhr University Bochum, including the work councils, the data protection officer, the communication department, and the IT department. Finally, we thank all participants for taking part in the study. This work was supported by the Sponsor Research Center Trustworthy Data Science and Security https://rc-trust.ai (https://rc-trust.ai), one of the Research Alliance Centers within the UA Ruhr (https://uaruhr.de).
@@ Appendix A Demographics
Figure A1. Percentage of participants using different chatbot systems.
Bar chart showing the percentage of participants for each usage category of the university’s customized LLMaaS chatbot and ChatGPT. The y-axis ranges from 0% to 70% in increments of 10%. From left to right on the x-axis: Both Systems – 22.1% (n = 116), Neither System – 23.4% (n = 123), Custom university LLMaaS Chatbot – 35.4% (n = 186), ChatGPT – 63.3% (n = 333), Only custom university LLMaaS Chatbot – 13.3% (n = 70), Only ChatGPT – 41.3% (n = 217). Each bar is colored differently for visual distinction.
@@ Appendix B Correlation Tables
Notes. $\alpha$ = Cronbach’s Alpha. Values in brackets = 95% CI. * $p<.05$, ** $p<.01$.
@@
Notes. $\alpha$ = Cronbach’s Alpha. Values in brackets = 95% CI. * $p<.05$, ** $p<.01$.
@@ Appendix C Descriptive Results
Figure C1. Usage frequency of the university’s customized LLMaaS chatbot vs. ChatGPT in the subsample of both systems users ($n=116$).
Grouped bar chart comparing usage frequency for the university’s customized LLMaaS chatbot and ChatGPT. The customized chatbot is represented by orange bars, and ChatGPT by blue bars. The y-axis is labeled “Percentage of Users (%)” and ranges from 0 to about 35 in regular increments. From left to right, the x-axis categories are: Once a month; Several times a month; Once a week; Several times a week; Once a day; Several times a day. Each bar is labeled with the count of users (n). Counts by category are: Once a month — customized chatbot n=15, ChatGPT n=18; Several times a month — customized chatbot n=38, ChatGPT n=29; Once a week — customized chatbot n=18, ChatGPT n=10; Several times a week — customized chatbot n=29, ChatGPT n=33; Once a day — customized chatbot n=5, ChatGPT n=4; Several times a day — customized chatbot n=11, ChatGPT n=22.
Figure C2. Purpose of use for the university’s customized LLMaaS chatbot vs. ChatGPT in the subsample of both systems users ($n=116$).
Grouped bar chart comparing purposes of use for the university’s customized LLMaaS chatbot and ChatGPT. The customized chatbot is represented by orange bars, and ChatGPT by blue bars. The y-axis is labeled “Percentage of Users (%)” and ranges from 0 to about 70 in increments of 10. From left to right, the x-axis categories are: Reviewing/Improving of Texts — customized chatbot n=82, ChatGPT n=82; Creation of Texts — customized chatbot n=63, ChatGPT n=71; Information Search — customized chatbot n=61, ChatGPT n=78; Programming/Code Review — customized chatbot n=55, ChatGPT n=51; Analyses/Summaries/Protocols — customized chatbot n=47, ChatGPT n=49; Translations — customized chatbot n=42, ChatGPT n=49; Creation of Images — customized chatbot n=28, ChatGPT n=33; Conversation Partner — customized chatbot n=15, ChatGPT n=28; Use in Seminars/Courses — customized chatbot n=13, ChatGPT n=14; Other — customized chatbot n=9, ChatGPT n=8.
@@ Appendix D Controlling for Confounding Variables
Table D1. Paired t-test comparisons of the customized LLMaaS chatbot (UniGPT) vs. ChatGPT – Non-switchers ($n=68$)
Privacy concerns (H5) 2.32 0.89 3.46 0.80 -9.34 67 ¡ 0.001 0.00833 -1.13 [-1.44, -0.82]
Perceived hallucinations (H4) 3.10 1.05 3.86 0.84 -6.41 67 ¡ 0.001 0.01 -0.78 [-1.05, -0.50]
Trust (H1) 3.60 0.71 3.09 0.81 5.97 67 ¡ 0.001 0.0125 0.72 [0.45, 0.99]
Sustainability-aware AI use (H6) 2.30 1.13 2.25 1.16 0.91 67 0.37 0.0167 0.11 [-0.13, 0.35]
Cautious behavior toward hallucinations (H3) 4.20 0.67 4.19 0.66 0.10 67 0.92 n/a4 0.01 [-0.23, 0.25]
Usefulness5 4.03 0.74 3.97 0.80 0.68 67 0.50 0.08 [-0.16, 0.32]
Ease of use5 4.52 0.67 4.46 0.58 1.02 67 0.31 0.12 [-0.11, 0.36]
Notes. 1 Holm-adjusted significance threshold per hypothesis rank (sequential correction). 2 Cohen’s d is reported as the standardized effect size for each comparison. 3 Confidence intervals (CI) for Cohen’s d are not adjusted for multiple comparisons. 4 Holm-adjusted thresholds are only assigned while p-values remain below $\alpha$; correction stops after the first non-significant test. 5 Ease of Use and Usefulness were analyzed exploratively and were not part of the Holm correction.
@@
PerceivedHallucinations: System $0.537^{***}$ $0.154$ 3.49-3.49 <0.001<0.001
PerceivedHallucinations: System × Knowledge -0.027 $0.046$ 0.60-0.60 $0.553$
@@
Appendix E Full-Collinearity Checks
Table E1. Full-collinearity VIFs for the mixed-effects model (System × Trust in University).
Predictor VIF
System 1.00
Trust in University (centered) 2.00
System × Trust in University 2.00
Max VIF 2.00
Table E2. Adjusted GVIFs for the ordinal regression model (System × Usage Frequency).
Predictor GVIF Df GVIF1/(2⋅Df)
System 7.12 1 2.67
Usage Frequency (polynomial contrasts) 1.21 5 1.02
System × Usage Frequency 8.31 5 1.24
Max adjusted GVIF 2.67
Appendix F Exploratory Analyses
[…]